good morning. My problem is this:
For each of the following functions, find a.f(4), b.f(-3), c.f(2.7, d. f(-4.9)
f(x) = -2x+4 if x less than or equals 1
3 if 1 < x <4
x+1 if x is more than or equal to 4
Can you help me with this problem? Than you for all of your help.

Found 2 solutions by hokies, vksarvepalli:
Answer by hokies(65)

f(4) = 4+1 = 5
f(-3) = -2*-3 + 4 = 6+4 = 10
f(2.7) = 3
f(-4.9) = -2*-4.9 + 4 = 9.8 + 4 = 13.8

Answer by vksarvepalli(154)

it is a single function,it is defined as such
a)f(4)
f(x) = -2x+4 if x less than or equals 1
3 if 1 < x <4
x+1 if x is more than or equal to 4
when x=4 f(x) is defined as x+1
so f(4)=5
b)f(-3)
-3 is less than 1
so f(x) = -2x+4
therefore f(-3)=(-2*-3)+4
=10
c)f(2.7)
1<2.7<4
therefore f(2.7)=3
d)f(-4.9)
-4.9<1
therefore f(-4.9)=(-2*-4.9)+4
=9.8+4
=13.8
f(-4.9)=13.8